Best Bicycles manufactures three different types of bikes: the Tiny Tike, the Adult Aero, and the Mountain Monger. Assume that Best Bicycles operates two shifts, each with 1,300 hours per year. Bike Demand Forecast (units/year) Processing Time (minutes/unit) Lot Size (# of bikes) Setup Time (minutes/setup) Tiny Tike 13,920 9.5 8 22 Adult Aero 89,470 40.8 46 44 Mountain Monger 17,290 34.0 14 56 Calculate the required capacity for this year’s production. Note that the times are given for assembly lines, so capacity calculations should be in terms of the number of lines necessary. (Round up your answer to the next whole number.)
Answer #1
Model Forecast Processing time in Min Lot size Set up time
Tiny Tike 13920 9.5 8 22
Adult Aero 89470 40.8 46 44
Mountain Monger 17290 34 14 56

Time per lot for Tiny tike = 9.5x8+22 = 98 min.

Number of lots that would be produced per year = 1300x2 x60 / 98 =1592 lots = 1592x8 =12736 bikes

NUmber of assembly lines needed for Tiny Tike = 13920/12736 = 1.093 = 2

Time per lot for adult aero = 40.8x46 +44 = 1920.8

No. of lots per year = 156000/1920.8 =81.216 lots = 3735 bikes

Number of assembly lines = 89470 /3735 = 23.95 =24

Time per lot of Mountain monger = 34x14+56 =532 min

No, of lots per year = 156000 /532 = 293.23

Number of bikes = 293.23x14 =4105 bikes

Number of assembly lines = 17290 /4105 = 4.21 =5